OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is even and 1 die shows a 4? A 6 X 6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

If 1 die shows a 4, what restrictions do we have on the other dice to ensure the sum is even?

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

Right, the other dice must show an even number. Which has a 50% chance. (3/6 = 1/2 = 50%) And what are the odds of rolling a 4 on one die? \[P(\text{A AND B}) = P(A) \times P(B)\]

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

Right. So if the first event has probability 1/6 and second event has probability 1/2 What is the probability to get A AND B ?

OpenStudy (anonymous):

1/12?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

If you buy a raffle ticket and there is only one prize available, you can double your probability of winning that prize by buying two tickets. true or false